OpenStudy (anonymous):

Solve the equation 7n – 2 = 5n + 10 for n.

OpenStudy (anonymous):

So move all the bits with n in them to one side and then move all the bits without n in them to the other side

OpenStudy (anonymous):

n = 6

OpenStudy (anonymous):

I don't want the answer I just want someone to show me how how to work it out please

OpenStudy (anonymous):

sure what happens if you add 2 to both sides of the equation

OpenStudy (anonymous):

like this \[7n -2 +2 = 5n +10 +2\]

OpenStudy (anonymous):

7n-2 = 5n +10 7n-5n -2 = 5n-5n +10 2n -2 = 10 2n -2+2=10+2 2n = 12 2n/2 = 12/2 n = 6

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

Yes this makes sense Thank's!